Subject: KAMI/KTeV EVE Shift Summary -- 1/18/2000 (Tue.)
From: Vivian O'Dell (firstname.lastname@example.org)
Date: Fri Feb 04 2000 - 15:31:28 CST
Shift Summary for EVE shift, 1/18/2000 (Tue.)
On Shift: Ron Ray and Earl Swallow + "the usual suspects ...."
Hours of beam available at all intensities: 8.2 (out of 8.5)
Hours of beam available at nominal intensity: 8.2
Hours of KAMI test data: 5.6
Hours of special runs (laser scans, muon runs...): 0.0
Accounting of lost time & intensity:
Accesses for necessary KAMI test run changes &
brief accelerator problems.
Runs ended this shift:
15592 -- KAMI test: LAMPPI.
15593 -- KAMI test: KAMI_KE3.
15595 -- KTEV_PED
15596 -- KAMI test: KAMI_MUON
Tapes brought to FCC this shift:
Accesses during shift:
Several -- see below.
Overview of shift:
KAMI test running in full progress.
Description of Shift:
15592 15:30 EVE Shift Begins with a large cast in control room.
Lambda ---> p pi data taking in progress (Run 15592).
Regenerator in use as lambda production target.
15592 End Run 15592 -- LAMPPI.
Great proton E/P peak! [See pages Logbook 82-83.]
Access to remove regenerator & AGEL work.
Shadow absorber removed remotely.
Move on to Ke3.
15593 16:05 Start Run 15593 -- KAMI_KE3 -- SEM = 4 E12.
(UCAL hole pion tag & BAG hole electron tag)
Still no beam after access -- magnet problem.
16:20 BEAM returns.
36 16:50 KFREND
RK checks UCAL timing -- looks good.
92 18:04 KFREND, spills 92-95.
RK realizes that FPEDCAL has been running and
resetting thresholds, which is then undone by KFREND.
FPEDCAL was NOT running during LAMPPI runs.
It has been running during 15593 -- Sasha L kills it.
19:30 Call MCR to turn off beam to begin cool-down.
RickF coming in to move beam plug, etc.
15593 End Run 15593.
20:55 NM2 access completed.
+ Beam plug moved from East to West beam.
+ Movable absorber removed.
+ Common absorber to position 1 (20.6 in.)
+ Concern: rising radiation level near plug in NM2.
21:10 NM4 access to put AGEL behind BAG.
21:15 NM4AN set to low field for muon run.
15595 Do Pedestal Run 15595 -- update pedestals.
21:20 Beam Returns briefly.
15596 21:25 Start Run 15596 -- KAMI_MUON.
Beam OFF - lost Meson primary.
21:30 Beam returns -- BAG work with muons -- SEM = 2 E12.
24:00 Shift ends with BAG muon work in progress.
KAMI Logbook I, pages 81-89.
Follow up Notes on Previous Shift
Items to Follow up on Next Shifts
-- =============================================================================== Earl C. Swallow email: email@example.com Theophil W. Mueller Professor (or) firstname.lastname@example.org and Chairperson Department of Physics Phone: (630)-617-3577 Elmhurst College FAX: (630)-617-3735 190 Prospect Avenue Elmhurst, IL 60126-3296 http://www.elmhurst.edu/~earls =============================================================================== "Books are lighthouses erected in the dark sea of time." -- from Gargoyles (The Scrolls of Merlin)
This archive was generated by hypermail 2b27 : Fri Feb 04 2000 - 15:31:33 CST